4 ms·
How?? The hole cards aren't enough - there are more cards coming on the table. Even if you know that your nothing beats his nothing you are risking a lot on a
by deanCommie 4y ago
How??
The hole cards aren't enough - there are more cards coming on the table.
Even if you know that your nothing beats his nothing you are risking a lot on a small advantage that his nothing won't turn into something.
- alasdair_ 4y agoThere is no more betting to go, so it does make sense. You are 47% to win but it’s +EV because there is already money in the pot so the pot odds favor you. That said, there are far better spots to do this and her other plays earlier were less than great.
- jstx1 4y ago> How?? After the shove on the turn the pot is $161k, she needs to call $109k. She needs 109/(109+161) = 40.3% equity to break even. So if she has more than 40.3% equity, she should call. If you plug in their cards on the turn into an equity calculator, you'll see that she has 54.55% equity (not accounting for dead cards like the stream calculation but either way it's the same outcome) so it's +EV for her to call.
- mod 4y agoIf Garrett turned his hand face up, she makes about 20k by calling on the turn.